# A Survey on the Safety and Security Threats of Computer-Using Agents: JARVIS or Ultron?

### v2

@@ 4 Taxonomy of Existing Defenses
4 Data Sanitization $\blacklozenge$ $\checkmark$ Ex. 4
@@
6 Output Monitoring $\blacklozenge$ $\checkmark$ In. 3 4 Ex. 6 7
7 Model Inspection $\blacklozenge$ $\checkmark$ Ex. 2 4
8 Cross-Verification $\blacklozenge$ $\checkmark$ $\checkmark$ Ex. 1 3
@@
13 Architecture Refinement $\blacklozenge$ $\checkmark$ $\checkmark$ In. 2 7 8 Ex. 5
@@ 4.2 Defense Categories
For example, research reveals how visual elements on mobile interfaces can be manipulated to trigger unintended behaviors in GUI agents (Yang et al., 2024c). As a defense, they suggest sandboxing agent execution within constrained environments that monitor for risky API calls, and filtering GUI event access to minimize potential injection vectors (Yang et al., 2024c; Zhang et al., 2023).
@@
Input Validation  Input validation is a security measure that involves verifying and sanitizing user inputs to prevent the system from processing malicious or unintended commands. This strategy is predominantly applied in single-agent models, focusing on scrutinizing prompts to ensure they do not contain harmful instructions or malicious injections. Within the agent framework, input validation operates primarily at the perception level, where the agent interprets and understands user inputs. The primary threat addressed by this method is jailbreak attacks, where adversaries craft inputs designed to bypass the model’s safety mechanisms and elicit unauthorized behaviors.
For example, AutoDroid uses a privacy filter to mask personal information before prompts are sent (Wen et al., 2023). A similar filter also exists in (Zhang et al., 2024c). Additionally, in (Kumar et al., 2024), researchers observed that LLM-based browser agents are trained with safeguards to refuse harmful instructions in chat settings. The study introduced the Browser Agent Red-teaming Toolkit (BrowserART), which comprises 100 diverse browser-related harmful behaviors.
@@
For example, in (Debenedetti et al., 2024), researchers introduced a structured evaluation environment to test and refine defensive prompting techniques. The study demonstrated that carefully crafted counter-prompts and reinforcement-based instruction tuning could significantly reduce the success rate of prompt injection attacks, enhancing model robustness (Debenedetti et al., 2024). Similarly, it was recommended that more detailed defensive prompts and robust content filtering should be used to enhance defense efficiency (Zhang et al., 2024c). Moreover, a safety prompt is introduced to instruct the agent to ignore malicious inconsistencies in  (Wu et al., 2024a). Also, experiments are done in (Chen et al., 2025b) to investigate the efficiency of this strategy.
@@
Data Sanitization  It refers to a process in machine learning that involves detecting and removing malicious or corrupted data from training datasets to ensure the integrity and security of models. Current discussion regarding this strategy mainly lies in the single-agent model, targeting at preventing malicious triggers during its reasoning and planning phase (Yang et al., 2024b). This preventive measure is essential to protect models from various attacks, such as backdoor and adversarial attacks.
@@
The primary focus of this method is the model component of the agent framework. By exposing models to adversarial examples during training, they learn to withstand such perturbations, thereby improving their robustness. This method specifically targets adversarial attacks, which involve subtle input modifications that can cause models to make incorrect predictions (Wu et al., 2024a).
For example, researchers demonstrated that Computer-Using Agents (CUAs) could be compromised through minimal perturbations to visual inputs, affecting their visual grounding (Wu et al., 2024a). By adversarial training, models can learn to recognize and resist these manipulations, thereby enhancing their task completion rate.
@@
Output Monitoring  It refers to a strategy that involves continuously observing and evaluating the outputs of language models to ensure they align with user intentions and do not produce undesired actions. This approach is primarily applied in single-agent systems, focusing on the model component within the action phase of the agent framework. It aims to address threats such as misalignment, where the agent’s actions diverge from user expectations, and hallucination, where the model generates incorrect or nonsensical information. Additionally, actions resulting in system sabotage or related to malicious usage, such as web hacking, could also be intercepted by this approach.
For instance, in the study (Fang et al., 2024a), the authors introduce InferAct, a novel approach that leverages the belief reasoning ability of large language models, grounded in Theory-of-Mind, to detect misaligned actions before execution. InferAct alerts users for timely correction, preventing adverse outcomes and enhancing the reliability of LLM agents’ decision-making processes (Fang et al., 2024a). Additionally, the Task Executor in AutoDroid verifies the security of an output action and asks the user to confirm if the action is potentially risky (Wen et al., 2023). Moreover, TrustAgent includes a post-planning inspection before tool calls (Hua et al., 2024).
@@
Model Inspection  This method detects malicious manipulations or compromised logic by examining internal model behaviors and parameters (Wang et al., 2025; Yang et al., 2024b). It applies to both single-agent and multi-agent systems, targeting the model component of the agent, and operates within the brain of the agent framework. Model inspection defends against critical threats such as backdoor attacks and prompt injection attacks by surfacing anomalous activity patterns or internal inconsistencies.
@@
It focuses on monitoring the behaviors of agents during inference or interaction to detect deviations from expected model outputs or communication topologies. It is especially relevant in multi-agent systems, where interactions can reveal inconsistencies in decision-making caused by compromised agents. For instance, a graph-based monitoring system was introduced to detect adversarially influenced agents by analyzing the topological communication patterns across agents (Wang et al., 2025). The system was able to isolate and prune suspect nodes based on anomaly scores derived from communication flows (Wang et al., 2025). Furthermore, a Graphormer model can analyze a dynamic spatio-semantic safety graph that captures both spatial and contextual risk factors in real-time to detect hazards (Huang et al., 2025).
@@
This involves inspecting the internal parameters of a trained model to identify hidden triggers or abnormal value distributions indicative of backdoor implantation. This approach is particularly relevant for single-agent systems. For example, the authors perform weight-based inspection of transformer layers to identify neurons with disproportionately high influence tied to specific trigger tokens in  (Yang et al., 2024b). The analysis revealed clear distinctions between clean and poisoned models, suggesting that weight-level scrutiny can expose embedded backdoors (Yang et al., 2024b). Additionally, (Zhu et al., 2025) proposed an automatic memory-audit step after every task, which flags anomalies in the agent’s internal memory traces to detect hidden backdoors.
@@
Cross Verification  This is a collaborative defense strategy in multi-agent systems where multiple agents independently process the same task or instruction and validate each other’s outputs to ensure consistency and correctness (Zeng et al., 2024). This method primarily targets the model component of the agent framework and operates across both the brain and action stages, with the aim of defending against jailbreak and adversarial attacks that may manipulate a single agent’s output to produce harmful or unauthorized behavior.
In the context of jailbreak prevention, cross-verification enables redundancy and consensus among agents, thereby reducing the likelihood that a single compromised response propagates through the system. For example, Zeng et al. propose a multi-agent defense architecture where a guard agent cross-validates the output of a task agent (Zeng et al., 2024). If the task agent generates potentially harmful content in response to a jailbreak attempt, the guard agent flags the behavior and halts execution, effectively mitigating the attack (Zeng et al., 2024). Additionally, AgentOccam uses a Judge agent to assess every candidate action and picks the one with the least risk (Yang et al., 2024a). Similarly, AGrail utilizes multiple checker agents to verify every candidate action before execution (Luo et al., 2025).
@@
Continuous Learning and Adaptation  Continuous learning and adaptation refers to the capability of agents to dynamically update their internal models based on new interactions, environments, or user feedback, thereby improving their long-term safety and robustness (Tian et al., 2023). This strategy is primarily discussed in the context of multi-agent systems, targeting the model as the primary defense component and the user as a secondary influence. Operating within the brain of the agent framework, this method aims to address prompt injection attacks by enabling agents to detect and adapt to adversarial prompts over time.
@@
It refers to the agent’s ability to autonomously adjust its reasoning or decision-making strategy based on past experiences and outcomes. LLM-based agents that re-encode their internal state across tasks are better at identifying unsafe instructions and suggest using performance memory or task replay buffers to evolve the agent’s policy over time (Tian et al., 2023; Luo et al., 2025). This helps reduce the success rate of prompt injection attacks by enabling agents to learn from near-miss or failed tasks.
@@
It leverages feedbacks from human users to realign the agent’s behavior with user expectations. In the same study, the authors show that agents assisted with user feedback—such as warning prompts or confirmations before execution—exhibited more cautious and aligned behavior when encountering ambiguous or adversarial inputs (Tian et al., 2023). This aligns with the idea proposed in (Ma et al., 2024) that human-in-the-loop designs enhance agent safety in real-world, evolving task environments.
@@
Transparentize  Transparentize refers to the implementation of mechanisms that enhance the transparency and interpretability of AI agents, thereby improving trust and safety in their operations. This strategy is particularly relevant in single-agent systems, focusing primarily on the model component and secondarily on the user component within the brain of the agent framework. It addresses threats such as hallucination—where the agent generates incorrect or nonsensical information—and misalignment, where the agent’s actions diverge from user intentions.
@@
This entails recording the actions and decisions made by AI agents to provide a traceable history of their operations. Maintaining detailed logs is recommended to monitor agent behavior, facilitate debugging, and ensure accountability (Sager et al., 2025).
@@
Topology-Guided  Topology-guided strategies enhance the security of multi-agent systems by analyzing and leveraging the structural relationships among agents to detect and mitigate adversarial threats (Wang et al., 2025). This approach is particularly relevant in multi-agent systems, focusing primarily on the model component within the brain and action phases of the agent framework. It addresses threats such as prompt injection attacks by examining the communication patterns and interactions among agents.
@@
It monitors the communication and interaction patterns among agents to identify anomalies that may indicate security breaches. For example, a multi-agent utterance graph could be constructed to monitor interactions and employ graph neural networks to detect anomalous communication flows that could signify prompt injection attacks (Wang et al., 2025).
@@
It focuses on designing the agent network topology to be robust against potential attacks. This includes strategies such as edge pruning, where connections to compromised agents are severed to prevent the spread of malicious information. The same study demonstrates that by adjusting the network topology through edge pruning, the system can effectively contain and mitigate the impact of detected attacks (Wang et al., 2025).
@@
For example, grounding inputs by combining element-attribute, textual-choice, and image-annotation cues dramatically reduces mis-click rates on web tasks (Zheng et al., 2024). Additionally, MobileFlow augments its pipeline with a hybrid visual encoder and Mixture of Experts (MoE) alignment training, boosting image interpretation on Android  (Nong et al., 2024). On the PC side, PC-Agent introduces an active perception module that uses A11y-tree parsing with OCR, achieving fine-grained element localisation in complex desktop windows (Liu et al., 2025). Finally, AgentOccam introduces observation-space alignment and page-simplification to address the excessive context length issue (Yang et al., 2024a).
@@
A representative approach is the chain-of-action prompt: it requires the agent to emit a full future-action plan before each execution step, cutting scheduling faults in half (Zhang and Zhang, 2023). Mobile-Bench extends this idea to multi-agent collaboration with a three-level (instruction, sub-task, action) hierarchy that decomposes long-horizon commands and reduces decision-making difficulties (Deng et al., 2024). AutoDroid lowers response latency by caching an LLM-generated guideline once per task, then delegating step-level binding to lightweight vision models (Wen et al., 2023). Complementarily, the PC-Agent framework allocates specialised Manager, Progress and Decision agents to refine and verify plans before execution, boosting success on 20-step desktop workflows (Liu et al., 2025).
@@
It refers to specific regulatory frameworks and industry standards pre-defined for CUAs to comply with. For example, a comprehensive benchmark (Zhang et al., 2024e) is introduced to assess the safety of large language model agents, ensuring they meet predefined safety standards and operate within acceptable risk parameters. Additionally, GameChat employs pre-defined Control Barrier Functions to define safe operational boundaries for each agent in a multi-agent system, ensuring agents’ trajectories remain within safe limits, preventing collisions  (Mahadevan et al., 2025). The game-theoretic strategy satisfying Subgame Perfect Equilibrium in GameChat further prevents agents from deviating from the agreed-upon strategies at any point, promoting consistent adherence to safe navigation protocols (Mahadevan et al., 2025). Moreover, ShieldAgent extracts verifiable rules from policy documents, structures them into a set of action-based probabilistic rule circuits, and associates specific agent actions with corresponding constraints (Chen et al., 2025c). Continuous verification ensures real-time standards adherence (Chen et al., 2025c).
@@
Assesses the robustness of multimodal web agents against adversarial attacks originating from the environment. 200 adversarial tasks Modifying open-source data Benign SR, ASR Rule
@@
Evaluates the safety of web agents by testing policy adherence and risk mitigation, focusing on external attacks and internal misalignments. 235 policy-enriched tasks Modifying open-source data CuP, Partial CuP Rule
@@
Assesses the safety of browser agents against harmful interactions, content, and jailbreak. 100 harmful browser-related behaviors Modifying open-source data ASR LLM
@@
(Chen et al., 2025c)
Tests agent safety against adversarial instructions and policy violations across web environments and risk categories. 960 web instructions, 3110 unsafe trajectories Modifying open-source data Accuracy, FPR, Recall, Inference Cost Rule
@@
General R-Judge
@@
Evaluates LLM agents’ safety awareness about multiple risks, with prompt injection attacks and complex environment challenges. 569 records of multi-turn agent interaction Modifying open-source data with ChatGPT F1 score, Recall, Specificity, Effectiveness Manual, LLM
@@
Evaluates agents’ safety regulations into planning across domains and risks. 144 data points Modifying open-source data Helpfulness, Safety, Total Correct Prefix, SSR LLM, Rule
@@
Evaluates LLM agents’ safety against jailbreaks and misalignments across risks. 2000 test cases with 10 failure modes and 349 environments Modifying open-source data Safety Score LLM, Rule
Table 4: An overview of computer-using agents (CUAs) safety benchmarks.

### v3

@@ 4.1 Defense Overview
4 Data Sanitization $\blacklozenge$ $\checkmark$ Ex. 4 5
@@
6 Output Monitoring $\blacklozenge$ $\checkmark$ In. 3 4 Ex. 7 8
7 Model Inspection $\blacklozenge$ $\checkmark$ Ex. 2 4 5
8 Cross-Verification $\blacklozenge$ $\checkmark$ $\checkmark$ Ex. 1 3 5
@@
13 Planning-Centric Architecture Refinement $\blacklozenge$ $\checkmark$ $\checkmark$ In. 2 7 8 Ex. 6
@@ 4.2 Defense Categories
For example, research reveals how visual elements on mobile interfaces can be manipulated to trigger unintended behaviors in GUI agents (Yang et al., 2024c). As a defense, they suggest sandboxing agent execution within constrained environments that monitor for risky API calls, and filtering GUI event access to minimize potential injection vectors (Yang et al., 2024c; Zhang et al., 2023). Additionally, GameChat uses control-barrier functions to constrain each agent’s trajectory to a safe region, preventing collisions and deadlocks in cluttered spaces (Mahadevan et al., 2025). Moreover, the framework in (Huang et al., 2025) builds a dynamic spatio-semantic safety graph that monitors real-time hazards and adaptively refines task plans to enforce safe execution.
@@
Input Validation  Input validation is a security measure that involves verifying and sanitizing user inputs to prevent the system from processing malicious or unintended commands (Ferrag et al., 2025). This strategy is predominantly applied in single-agent models, focusing on scrutinizing prompts to ensure they do not contain harmful instructions or malicious injections. Within the agent framework, input validation operates primarily at the perception level, where the agent interprets and understands user inputs. The primary threat addressed by this method is jailbreak attacks, where adversaries craft inputs designed to bypass the model’s safety mechanisms and elicit unauthorized behaviors.
For example, AutoDroid uses a privacy filter to mask personal information before prompts are sent (Wen et al., 2023). A similar filter also exists in (Zhang et al., 2024c). Additionally, in (Kumar et al., 2024), researchers observed that LLM-based browser agents are trained with safeguards to refuse harmful instructions in chat settings. The study introduced the Browser Agent Red-teaming Toolkit (BrowserART), which comprises 100 diverse browser-related harmful behaviors. Moreover, the authors in (Tshimula et al., 2024) apply pattern matching and high-precision filters to incoming prompts to detect and strip out jailbreak payloads before they reach the LLM. PromptArmor runs a lightweight LLM pre-processor that scans and sanitizes user inputs, removing any suspicious sub-prompts before they’re forwarded to the agent (Shi et al., 2025a). Also, RTBAS employs dynamic information-flow control and dual dependency screeners to vet tool calls, automatically ensuring confidentiality and integrity without constant user confirmation (Zhong et al., 2025).
@@
For example, in (Debenedetti et al., 2024), researchers introduced a structured evaluation environment to test and refine defensive prompting techniques. The study demonstrated that carefully crafted counter-prompts and reinforcement-based instruction tuning could significantly reduce the success rate of prompt injection attacks, enhancing model robustness (Debenedetti et al., 2024). Similarly, it was recommended that more detailed defensive prompts and robust content filtering should be used to enhance defense efficiency (Zhang et al., 2024c). Moreover, a safety prompt is introduced to instruct the agent to ignore malicious inconsistencies in  (Wu et al., 2024a). Also, experiments are done in (Chen et al., 2025d) to investigate the efficiency of this strategy.
@@
Data Sanitization  It refers to a process in machine learning that involves detecting and removing malicious or corrupted data from training datasets to ensure the integrity and security of models. Current discussion regarding this strategy mainly lies in the single-agent model, targeting at preventing malicious triggers during its reasoning and planning phase (Yang et al., 2024b; Jones et al., 2025; Wang et al., 2025b). This preventive measure is essential to protect models from various attacks, such as backdoor and memory injection attacks.
@@
The primary focus of this method is the model component of the agent framework. By exposing models to adversarial examples during training, they learn to withstand such perturbations, thereby improving their robustness (Yu et al., 2025). This method specifically targets adversarial attacks, which involve subtle input modifications that can cause models to make incorrect predictions (Wu et al., 2024a).
For example, researchers demonstrated that Computer-Using Agents (CUAs) could be compromised through minimal perturbations to visual inputs, affecting their visual grounding (Wu et al., 2024a; Yu et al., 2020). By adversarial training, models can learn to recognize and resist these manipulations, thereby enhancing their task completion rate, as demonstrated in AutoSafe, which synthesizes diverse risk scenarios and uses them as on-the-fly adversarial examples during fine-tuning to markedly improve agent robustness (Zhou et al., 2025).
@@
Output Monitoring  It refers to a strategy that involves continuously observing and evaluating the outputs of language models to ensure they align with user intentions and do not produce undesired actions (Shi et al., 2025b). This approach is primarily applied in single-agent systems, focusing on the model component within the action phase of the agent framework. It aims to address threats such as misalignment, where the agent’s actions diverge from user expectations, and hallucination, where the model generates incorrect or nonsensical information. Additionally, actions resulting in system sabotage or related to malicious usage, such as web hacking, could also be intercepted by this approach.
For instance, in the study (Fang et al., 2024a), the authors introduce InferAct, a novel approach that leverages the belief reasoning ability of large language models, grounded in Theory-of-Mind, to detect misaligned actions before execution. InferAct alerts users for timely correction, preventing adverse outcomes and enhancing the reliability of LLM agents’ decision-making processes (Fang et al., 2024a). Additionally, the Task Executor in AutoDroid verifies the security of an output action and asks the user to confirm if the action is potentially risky (Wen et al., 2023). Moreover, TrustAgent includes a post-planning inspection before tool calls (Hua et al., 2024). VeriSafe Agent auto-formalizes user instructions into a DSL specification and checks each GUI operation at runtime, blocking any action that fails logic checks (Lee et al., 2025).
@@
Model Inspection  This method detects malicious manipulations or compromised logic by examining internal model behaviors and parameters (Wang et al., 2025d; Yang et al., 2024b). It applies to both single-agent and multi-agent systems, targeting the model component of the agent, and operates within the brain of the agent framework. Model inspection defends against critical threats such as backdoor attacks, prompt injection attacks, and memory injection attacks by surfacing anomalous activity patterns or internal inconsistencies.
@@
It focuses on monitoring the behaviors of agents during inference or interaction to detect deviations from expected model outputs or communication topologies. It is especially relevant in multi-agent systems, where interactions can reveal inconsistencies in decision-making caused by compromised agents. For instance, a graph-based monitoring system was introduced to detect adversarially influenced agents by analyzing the topological communication patterns across agents (Wang et al., 2025d). The system was able to isolate and prune suspect nodes based on anomaly scores derived from communication flows (Wang et al., 2025d). Furthermore, a Graphormer model can analyze a dynamic spatio-semantic safety graph that captures both spatial and contextual risk factors in real-time to detect hazards (Huang et al., 2025).
@@
This involves inspecting the internal parameters of a trained model to identify hidden triggers or abnormal value distributions indicative of backdoor implantation. This approach is particularly relevant for single-agent systems. For example, the authors perform weight-based inspection of transformer layers to identify neurons with disproportionately high influence tied to specific trigger tokens in  (Yang et al., 2024b). The analysis revealed clear distinctions between clean and poisoned models, suggesting that weight-level scrutiny can expose embedded backdoors (Yang et al., 2024b). Additionally, (Zhu et al., 2025b) proposed an automatic memory-audit step after every task, which flags anomalies in the agent’s internal memory traces to detect hidden backdoors.
@@
Cross Verification  This is a collaborative defense strategy in multi-agent systems where multiple agents independently process the same task or instruction and validate each other’s outputs to ensure consistency and correctness (Zeng et al., 2024). This method primarily targets the model component of the agent framework and operates across both the brain and action stages, with the aim of defending against jailbreak, adversarial attacks, and backdoor attacks that may manipulate a single agent’s output to produce harmful or unauthorized behavior.
In the context of jailbreak prevention, cross-verification enables redundancy and consensus among agents, thereby reducing the likelihood that a single compromised response propagates through the system. For example, Zeng et al. and Huang et al. propose a multi-agent defense architecture where a guard or review agent cross-validates the output of a task agent (Zeng et al., 2024; Huang et al., 2024). If the task agent generates potentially harmful content in response to a jailbreak attempt, the guard agent flags the behavior and halts execution, effectively mitigating the attack (Zeng et al., 2024). Additionally, AgentOccam uses a Judge agent to assess every candidate action and picks the one with the least risk (Yang et al., 2024a). Similarly, GuardAgent spins up a separate guardian LLM that re-evaluates the primary agent’s outputs against knowledge bases, vetoing any unsafe recommendations (Xiang et al., 2024). Moreover, AGrail utilizes multiple checker agents to verify every candidate action before execution (Luo et al., 2025b). Also, the approach in (Barua et al., 2025) uses multiple independent runs of the same prompt across agents and uses majority consensus to filter out jailbreak attempts. MELON executes each prompt twice, once normally and once with a masked injection, to compare outputs and flag any inconsistencies as injected content (Zhu et al., 2025a). For backdoor attacks, ReAgent performs dual-level consistency checks between planning thoughts and executed actions to detect and abort backdoor-triggered behaviors at inference time (Li et al., 2025); and PeerGuard leverages mutual reasoning among agents to cross-verify each other’s outputs and isolate poisoned agents in a multi-agent backdoor defense (Fan and Li, 2025).
@@
Continuous Learning and Adaptation  Continuous learning and adaptation refers to the capability of agents to dynamically update their internal models based on new interactions, environments, or user feedback, thereby improving their long-term safety and robustness (Tian et al., 2023; Zhan et al., 2025). This strategy is primarily discussed in the context of multi-agent systems, targeting the model as the primary defense component and the user as a secondary influence. Operating within the brain of the agent framework, this method aims to address prompt injection attacks by enabling agents to detect and adapt to adversarial prompts over time.
@@
It refers to the agent’s ability to autonomously adjust its reasoning or decision-making strategy based on past experiences and outcomes. LLM-based agents that re-encode their internal state across tasks are better at identifying unsafe instructions and suggest using performance memory or task replay buffers to evolve the agent’s policy over time (Tian et al., 2023; Luo et al., 2025b). This helps reduce the success rate of prompt injection attacks by enabling agents to learn from near-miss or failed tasks.
@@
It leverages feedback from human users to realign the agent’s behavior with user expectations. In the same study, the authors show that agents assisted with user feedback—such as warning prompts or confirmations before execution—exhibited more cautious and aligned behavior when encountering ambiguous or adversarial inputs (Tian et al., 2023). This aligns with the idea proposed in (Ma et al., 2024) that human-in-the-loop designs enhance agent safety in real-world, evolving task environments. For example, the study in (Zhang et al., 2025b) highlights user-initiated oversight mechanisms, such as manual correction loops and adaptive interface adjustments, enabling agents to learn from unintended outcome feedback and improve future interactions.
@@
Transparentize  Transparentize refers to the implementation of mechanisms that enhance the transparency and interpretability of AI agents, thereby improving trust and safety in their operations. This strategy is particularly relevant in single-agent systems, focusing primarily on the model component and secondarily on the user component within the brain of the agent framework. It addresses threats such as hallucination—where the agent generates incorrect or nonsensical information—and misalignment, where the agent’s actions diverge from user intentions to risky operations.
@@
This entails recording the actions and decisions made by AI agents to provide a traceable history of their operations. Maintaining detailed logs is recommended to monitor agent behavior, facilitate debugging, and ensure accountability (Sager et al., 2025). For example, the authors in (Chen et al., 2025b) propose in-context consent dialogues and user-facing risk indicators to increase transparency of GUI agent operations and empower users to make informed decisions.
@@
Topology-Guided  Topology-guided strategies enhance the security of multi-agent systems by analyzing and leveraging the structural relationships among agents to detect and mitigate adversarial threats (Wang et al., 2025d). This approach is particularly relevant in multi-agent systems, focusing primarily on the model component within the brain and action phases of the agent framework. It addresses threats such as prompt injection attacks by examining the communication patterns and interactions among agents.
@@
It monitors the communication and interaction patterns among agents to identify anomalies that may indicate security breaches. For example, a multi-agent utterance graph could be constructed to monitor interactions and employ graph neural networks to detect anomalous communication flows that could signify prompt injection attacks (Wang et al., 2025d).
@@
It focuses on designing the agent network topology to be robust against potential attacks. This includes strategies such as edge pruning, where connections to compromised agents are severed to prevent the spread of malicious information. The same study demonstrates that by adjusting the network topology through edge pruning, the system can effectively contain and mitigate the impact of detected attacks (Wang et al., 2025d).
@@
For example, grounding inputs by combining element-attribute, textual-choice, and image-annotation cues dramatically reduces mis-click rates on web tasks (Zheng et al., 2024). Additionally, MobileFlow augments its pipeline with a hybrid visual encoder and Mixture of Experts (MoE) alignment training, boosting image interpretation on Android  (Nong et al., 2024). On the PC side, PC-Agent introduces an active perception module that uses A11y-tree parsing with OCR, achieving fine-grained element localisation in complex desktop windows (Liu et al., 2025b). Finally, AgentOccam introduces observation-space alignment and page-simplification to address the excessive context length issue (Yang et al., 2024a).
@@
A representative approach is the chain-of-action prompt: it requires the agent to emit a full future-action plan before each execution step, cutting scheduling faults in half (Zhang and Zhang, 2023). Mobile-Bench extends this idea to multi-agent collaboration with a three-level (instruction, sub-task, action) hierarchy that decomposes long-horizon commands and reduces decision-making difficulties (Deng et al., 2024). AutoDroid lowers response latency by caching an LLM-generated guideline once per task, then delegating step-level binding to lightweight vision models (Wen et al., 2023). Complementarily, the PC-Agent framework allocates specialised Manager, Progress and Decision agents to refine and verify plans before execution, boosting success on 20-step desktop workflows (Liu et al., 2025b).
@@
It refers to specific regulatory frameworks and industry standards pre-defined for CUAs to comply with. For example, a comprehensive benchmark (Zhang et al., 2024e) is introduced to assess the safety of large language model agents, ensuring they meet predefined safety standards and operate within acceptable risk parameters. Additionally, GameChat employs pre-defined Control Barrier Functions to define safe operational boundaries for each agent in a multi-agent system, ensuring agents’ trajectories remain within safe limits, preventing collisions  (Mahadevan et al., 2025). The game-theoretic strategy satisfying Subgame Perfect Equilibrium in GameChat further prevents agents from deviating from the agreed-upon strategies at any point, promoting consistent adherence to safe navigation protocols (Mahadevan et al., 2025). Moreover, ShieldAgent extracts verifiable rules from policy documents, structures them into a set of action-based probabilistic rule circuits, and associates specific agent actions with corresponding constraints (Chen et al., 2025e). Continuous verification ensures real-time standards adherence (Chen et al., 2025e). Also, AgentSandbox operationalizes security principles like defense-in-depth and least privilege within agent lifecycles, embedding policy enforcement checkpoints that uphold confidentiality and integrity requirements (Zhang et al., 2025a).
@@
Assesses the robustness of multimodal web agents against adversarial attacks originating from the environment. 200 adversarial tasks Open-source data modification Benign SR, ASR Rule
@@
Evaluates the safety of web agents by testing policy adherence and risk mitigation, focusing on external attacks and internal misalignments. 235 policy-enriched tasks Open-source data modification CuP, Partial CuP Rule
@@
Assesses the safety of browser agents against harmful interactions, content, and jailbreak. 100 harmful browser-related behaviors Open-source data modification ASR LLM
@@
SafeArena
(Tur et al., 2025)
Evaluate deliberate misuse of autonomous web agents and introduces the ARIA risk framework. 250 safe and 250 harmful tasks Human curation with LLM assistance, Open-source data augmentation TCR, RR, Normalized Safety Score Rule, LLM, Manual
AgentDAM
(Zharmagambetov et al., 2025)
Measures inadvertent leakage of sensitive information by AI agents during web task execution. 246 tasks Human curation, Open-source utilization Utility, LR Rule, LLM
@@
(Chen et al., 2025e)
Tests agent safety against adversarial instructions and policy violations across web environments and risk categories. 960 web instructions, 3110 unsafe trajectories Open-source data modification Accuracy, FPR, Recall, Inference Cost Rule
WASP
(Evtimov et al., 2025)
Shows that even top-tier AI models can be deceived by simple, low-effort human-written injections in very realistic scenarios. 84 tasks Human curation TSR, Intermediate ASR Rule, LLM
VPI-Bench
(Cao et al., 2025)
Evaluates the robustness of CUAs and Browser-use agents to visual prompt injection across five popular web platforms. 306 test cases Human curation AR, ASR LLM
@@
Hijacking Jarvis
(Liu et al., 2025a)
Evaluates mobile GUI agents’ safety under unprivileged third-party UI manipulations by the AgentHazard framework. 3000+ attack scenarios Human creation, annnotation TSR, MR, $ACC_{safe}$, $ACC_{attack}$ Rule
Table 4: An overview of web and mobile based computer-using agents (CUAs) safety benchmarks.
Platform Benchmark Highlight Data Size Collection Metric Measure
General ToolEmu
(Ruan et al., 2023)
Evaluates safety failures of LM agents across diverse tool-driven scenarios. 36 toolkits, 144 test cases Human curation with LLM assistance Safety, Helpfulness LLM, Manual
R-Judge
@@
Evaluates LLM agents’ safety awareness about multiple risks, with prompt injection attacks and complex environment challenges. 569 records of multi-turn agent interaction Open-source data modification with ChatGPT F1 score, Recall, Specificity, Effectiveness Manual, LLM
@@
Evaluates agents’ safety regulations into planning across domains and risks. 144 data points Open-source data modification Helpfulness, Safety, Total Correct Prefix, SSR LLM, Rule
@@
HAICOSYSTEM
(Zhou et al., 2024)
Simulates multi-turn human–agent tool interactions to probe multi-dimensional safety risks. 132 scenarios, 8K simulated episodes Human creation, Open-source inspiration TARG, SYST, CONT, SOC, LEGAL, EFF, GOAL LLM
@@
Evaluates LLM agents’ safety against jailbreaks and misalignments across risks. 2000 test cases with 10 failure modes and 349 environments Open-source data modification Safety Score LLM, Rule
RedTeamCUA
(Liao et al., 2025)
Demonstrates that indirect prompt injection presents tangible risks for even advanced CUAs despite their capabilities and safeguards. 216 adversarial scenarios Human Curation SR, ASR, AR Rule, LLM
RiOSWorld
(Yang et al., 2025b)
Measures the risk intent and completion of MLLM-based agents during real-world computer manipulations. 492 risky tasks Human, Open-source, LLM RGC, RGI Rule, LLM
MLA-Trust
(Yang et al., 2025c)
Measures agent trustworthiness by orchestrating high-risk, interactive tasks, especially in multi-step interactions. 34 tasks Human creation, Open-source data augmentation Accuracy, MisRate, ASR, TS, RtE Rule LLM
GUI-Robust
(Yang et al., 2025a)
Reveal GUI agents’ substantial performance degradation in abnormal scenarios. 5318 tasks Semi-automated dataset construction paradigm Action &\& Coordinate Accuracy, TSR Rule
OS-Harm
(Kuntz et al., 2025)
Measures CUA safety across three harm types—deliberate user misuse, prompt injection, and model misbehavior. 150 tasks Human creation with LLM assistance, Open-source data augmentation Unsafe, TSR LLM
RAS-Eval
(Fu et al., 2025)
Evaluates security of LLM-based agents across simulated and real-world tool executions in diverse formats. 80 test cases, 3802 attack tasks Human collection, implementation TCR, TIR, TFR, score, ASR Rule
OpenAgent-Safety
(Vijayvargiya et al., 2025)
Evaluates agent safety when interacting with real tools across mixed environments including web and OS. 350 multi-turn, multi-user tasks Human curation with LLM assistance Unsafe Behavior Rates, Failure Rate, Disagreement Rate Rule, LLM
Table 5: An overview of general-purpose computer-using agents (CUAs) safety benchmarks.